# Clinically inspired enhance Explainability and Interpretability of an AI-Tool for BCC diagnosis based on expert annotation

### v2

@@ 2.3 Clinically-inspired Visual XAI
This section refers to the AI model at the top of Novel telemedicine diagnosis in Fig. 1 (BCC/Non-BCC classification). To develop this concept, expert-generated segmentations are used. As detailed in Sect. 2.1, a subset of samples was segmented by a specialist. A dermatologist performed an individual segmentation of each BCC pattern for each image (Fig. 3).
Understanding the decision-making process of convolutional neural networks (CNNs) is critical for clinical applications where interpretability is as important as performance metrics. The most common visualization techniques in XAI are activation maps and gradient-weighted class activation mapping (Grad-CAM). Activation Maps, as described by Zhou et al. [19], extract features from different network layers to show which patterns the model focuses on, although they may not always directly correlate with specific outcomes. On the other hand, Grad-CAM, as described in detail by Selvaraju et al. [10], uses the gradients of the predicted class that flow into the final convolutional layer to produce a localization map. The clinically-inspired explanation that we proposed in this paper is based on Grad-CAM.
@@
Figure 3: Example of the specialist segmentation job.
3 Results
3.1 Implementation details
Optimisation of the model as a key element of this project. Several critical hyperparameters were carefully adjusted in order to improve performance. The network architecture was trained in two stages. First, ImageNet’s transfer learning stage used the new AdamW schedule-free optimiser [20] with a learning rate of $1e^{-5}$ for 100 epochs, employing a mini-batch size of 32. A dropout rate of $0.3$ was established to serve as a regularisation technique to prevent overfitting.
As demonstrated in Sect. 2.1, class imbalance is a significant issue, especially in the multi-label classification of BCC patterns such as ML, MG, or SW. To address this problem, the Focal loss [21] was implemented, which is particularly useful when the distribution of samples is not uniform.
Due to the restricted scope of our database, we have employed a stratified k-fold cross-validation to guarantee a thorough assessment of the model’s efficacy. Specifically, a $5$-fold stratification [22, 23] was used, ensuring that each fold was well-representative of the whole dataset, maintaining the proportion of each class in every fold. This approach mitigates the risk of biased training and testing distributions, which is crucial in datasets with imbalanced class distributions.
For the classification model, we chose MobileNet-V2222, a lightweight deep neural network known for its efficiency and effectiveness in image classification tasks. To enhance the robustness of the model and account for variabilities in real-world dermoscopic imaging, several data augmentation techniques were applied [24, 25, 26]. These included rotation to mimic variations in the orientation of lesions in clinical images, random perspective to simulate different angles and perspectives from which images can be taken during examinations, and Gaussian blur to represent common issues of slight blur due to patient movement or focus inaccuracies in dermoscopic devices. These augmentation strategies not only help prevent overfitting, but also enhance the model’s ability to generalise to new, unseen images by simulating a variety of imaging conditions in real-world scenarios.
3.2 Clinically-inspired XAI: BCC diagnosis with additional label information
In general, the diagnostic performance of BCC / non-BCC is high, around $0.9$ for all metrics. However, the BCC pattern detection performance needs to be analysed with a deeper insight. Minority classes tend to achieve low recall because the AI tool trained with unbalanced databases tends to favour majority classes. As shown in Sect. 2.1, SW, MG and ML are underrepresented classes. Strategies such as data augmentation and advanced sampling, a one-vs-all strategy combined with stratified k-fold cross-validation helped to achieve a more balanced classification across patterns, thereby improving overall model performance. However, the metrics achieved should not be analysed in the same way as BCC/non-BCC performance. They should only be evaluated to the extent that they provide a correct explanation for binary classification. For the specialist, a tool that perfectly classifies all BCC patterns within a lesion is not particularly useful, as identifying even a single pattern can already indicate a positive BCC diagnosis. Therefore, the third section of Table 3 provides information relevant to the clinician, highlighting whether there is a BCC pattern detected by the specialist is present. Furthermore, this information is further supported by Table 4, which, as will be discussed later, shows whether the classification of each pattern is being extracted from the same region as identified by the dermatologist. This additional evaluation is summarised in the third section of Table 3, titled “Clinically-inspired XAI”. The table demonstrates that 73% of non-BCC lesions without any BCC patterns, 95% of non-BCC lesions with PN, and 99% of BCC lesions with at least one BCC pattern are correctly labelled as such.
BCC/Non-BCC
0.89 0.89 0.90 0.90
Pattern detection
OvidNests 0.65 ($2.48\cdot 10^{-3}$) 0.84 ($1.52\cdot 10^{-4}$) 0.53 ($3.78\cdot 10^{-5}$) 0.84 ($1.20\cdot 10^{-5}$)
Clinically-inspired XAI
All 0’s - - - 0.73
Pigment Network 0.94 0.96 0.97 0.95
BCC pattern detection 0.84 0.88 0.71 0.99
3.3 Clinically-inspired Visual XAI
This section aims to quantify the accuracy of the AI tool in focusing on the correct part of the lesion, specifically the BCC dermoscopic patterns identified by clinicians. To this end, the areas of the BCC pattern delineated by dermatologists will be compared with the activated areas of the model. This will provide a quantitative measure of the model’s agreement with human diagnostic criteria and demonstrate its ability to accurately identify critical features of BCC lesions.
(b) GradCAM heatmap showcasing areas of highest model activation.
(c) Integration of Grad-CAM heatmap and specialist segmentation.
Figure 4: Demonstration of various interpretative visualizations used in skin lesion analysis, combining expert segmentation and Grad-CAM heatmaps to aid in understanding the model’s decision-making process. This example is from a correct prediction sample.
Table 4: Statistics derived from estimation of conditional probability density functions of Grad-CAM within and outside the region of clinical interest.
Prediction Intersection Mean $P\left(z\left(x,y\right)\mid\text{Fg}\right)$ Mean $P\left(z\left(x,y\right)\mid\text{Bg}\right)$ Std $P\left(z\left(x,y\right)\mid\text{Fg}\right)$ Std $P\left(z\left(x,y\right)\mid\text{Bg}\right)$
Correct 0.24 0.57 0.16 0.14 0.22
Incorrect 0.32 0.33 0.14 0.01 0.21
@@ 4 Discussion
5 Acknowledgements
This work was supported by the Andalusian Regional Government (PROYEXCEL_00889) and is part of the project of the aid PID2021-127871OB-I00, funded by MCIN/AEI/10.13039/501100011033 and ERDF/European Union (NextGenerationEU/PRTR).
Conflict of Interest Statement
The authors declare that they have no conflict of interest regarding this study.

### v3

@@ 2.3 Multi-task learning (MTL) System: BCC Diagnosis with XAI-guided Pattern Recognition
2.4 Implementation details
Optimising the model was a key component of this work, with a particular focus on enhancing both classification performance and clinical interpretability. To this end, several critical hyperparameters were tuned through empirical experimentation. The training was conducted in two stages. In the first stage, transfer learning from ImageNet was employed using the AdamW schedule-free optimiser [defazio2024road_schedulefree], with a learning rate of $1\times 10^{-5}$ over 100 epochs and a mini-batch size of 32. A dropout rate of 0.3 was selected to regularise the training and reduce the risk of overfitting.
As demonstrated in Sect. 2.1, class imbalance is a significant issue, especially in the multi-label classification of BCC patterns such as ML, MG, or SW. To address this problem, the Focal loss [lin2018focal_loss] was implemented, which is particularly useful when the distribution of samples is not uniform.
Due to the restricted scope of our database, we have employed a stratified k-fold cross-validation to guarantee a thorough assessment of the model’s efficacy. Specifically, a $5$-fold stratification [hastie2009elements, Stone-KFold] was used, ensuring that each fold was well-representative of the whole dataset, maintaining the proportion of each class in every fold. This approach mitigates the risk of biased training and testing distributions, which is crucial in datasets with imbalanced class distributions.
For the classification model, we chose MobileNet-V22, a lightweight deep neural network known for its efficiency and effectiveness in image classification tasks. To enhance the robustness of the model and account for variabilities in real-world dermoscopic imaging, several data augmentation techniques were applied [AlexNet, cireşan2012multicolumn, simard2003best]. These included rotation to mimic variations in lesion orientation across clinical images, random perspective to simulate different acquisition angles during examination, and Gaussian blur to reflect common issues such as slight motion blur or focus inaccuracies in dermoscopic imaging. These augmentation strategies not only help prevent overfitting, but also enhance the model’s ability to generalise to new, unseen images by simulating a variety of imaging conditions in real-world scenarios.
3 Results
3.1 Pattern-Based Clinical Reasoning
In general, the diagnostic performance of BCC / non-BCC is high, around $0.9$ for all metrics. However, the BCC pattern detection performance needs to be analysed with a deeper insight. Minority classes tend to achieve low recall because the AI tool trained with unbalanced databases tends to favour majority classes. As shown in Sect. 2.1, SW, MG and ML are underrepresented classes. Strategies such as data augmentation and advanced sampling, a one-vs-all strategy combined with stratified k-fold cross-validation helped to achieve a more balanced classification across patterns, thereby improving overall model performance. However, the metrics achieved should not be analysed in the same way as BCC/non-BCC performance. They should only be evaluated to the extent that they provide a correct explanation for binary classification. For the specialist, a tool that perfectly classifies all BCC patterns within a lesion is not particularly useful, as identifying even a single pattern can already indicate a positive BCC diagnosis. Therefore, the third section of Table 3 provides information relevant to the clinician by assessing the model’s ability to detect at least one of the BCC patterns annotated by specialists, which is sufficient for a positive BCC diagnosis according to clinical criteria. Furthermore, this information is further supported by Table 4, which, as will be discussed later, shows whether the classification of each pattern is being extracted from the same region as identified by the dermatologist. This additional evaluation is summarised in the third section of Table 3, titled “Clinically-inspired XAI”. The table demonstrates that 73% of non-BCC lesions without any BCC patterns, 95% of non-BCC lesions with PN, and 99% of BCC lesions with at least one BCC pattern are correctly labelled as such.
BCC/Non-BCC
0.89 0.89 0.90 0.90
Pattern detection
Ovoid Nests 0.65 ($2.48\cdot 10^{-3}$) 0.84 ($1.52\cdot 10^{-4}$) 0.53 ($3.78\cdot 10^{-5}$) 0.84 ($1.20\cdot 10^{-5}$)
Clinically-inspired XAI
No-pattern (all zeros) 0.73
Pigment Network 0.94 0.96 0.97 0.95
BCC pattern detection 0.84 0.88 0.71 0.99
@@ 3.2 Clinically-inspired Visual XAI
3.2.1 Clinical Validation Framework
This section refers to the AI model at the top of Novel telemedicine diagnosis in Fig. 1 (BCC/Non-BCC classification). From a clinical perspective, the utility of an XAI system extends beyond simple visualization of model attention—it must demonstrate that the model’s focus aligns with the diagnostic criteria used by dermatologists. To address this requirement, we quantify the accuracy of the AI tool in focusing on the correct part of the lesion, specifically the BCC dermoscopic patterns identified by clinicians. To this end, the areas of the BCC pattern delineated by dermatologists are compared with the activated areas of the model, providing a quantitative measure of the model’s agreement with human diagnostic criteria and demonstrating its ability to accurately identify critical features of BCC lesions.
To develop this concept, expert-generated segmentations are used. As detailed in Sect. 2.1, a subset of 334 samples was segmented by a specialist. A dermatologist performed an individual segmentation of each BCC pattern for each image (Fig. 4). These expert annotations serve as the clinical ground truth against which the model’s attention mechanisms are validated, ensuring that XAI methods are not only technically robust but also tailored to clinical reasoning processes.
3.2.2 Expert-Model Agreement Assessment
To complement visual inspection, we quantify the spatial agreement between Grad-CAM activations and expert segmentations using conditional probability density functions, as described below.
To complement the descriptive analysis of Grad-CAM conditional distributions for correctly and incorrectly classified cases (upper part of Table 4), we performed an image-wise statistical comparison between activations inside and outside the dermatologists’ masks in the energy domain $z^{2}(x,y)$. As reported in the lower part of Table 4, statistically significant differences between foreground and background energy were observed in 328/334 images (98.2 %) at $\alpha=0.05$, with 322/334 images (96.5 %) remaining significant after Bonferroni correction and 328/334 images (98.2 %) after FDR (Benjamini–Hochberg) correction. The corresponding mean Cohen’s $d$ of 1.972 indicates a very large effect size, confirming that Grad-CAM concentrates substantially more energy within the region of clinical interest than in the surrounding background and providing strong statistical evidence for the foreground–background activation contrast.
(b) GradCAM heatmap showcasing areas of highest model activation.
(c) Integration of Grad-CAM heatmap and specialist segmentation.
Figure 3: Multi-panel visualization demonstrating the clinical validation of Grad-CAM activations against expert segmentations. (a) displays the original dermoscopic image with the expert segmentation boundary (red contour) delineating the clinically relevant lesion region; (b) shows the corresponding Grad-CAM heatmap indicating regions of highest model activation (intensity scale from low to high); (c) overlays the dermatologist’s manual segmentation on the Grad-CAM heatmap to assess spatial alignment; panel (d) presents the conditional probability density functions comparing Grad-CAM activation distributions inside (blue curve) versus outside (orange curve) the clinically relevant regions. The clear separation between distributions demonstrates that the model’s attention is appropriately focused on dermoscopically significant areas identified by clinical experts. This example illustrates a correctly predicted BCC case where the model’s decision-making aligns with clinical diagnostic criteria.
Table 4: Summary of Grad-CAM descriptive statistics and energy-based validation inside and outside the region of clinical interest.The upper section reports conditional means and standard deviations of foreground and background activations. The lower section presents the statistical comparison of integrated energy ($\sum z^{2}$) between both regions, including the proportion of significant $z$-tests with multiple-comparison corrections and the corresponding mean effect size (Cohen’s $d$), quantifying the strength of the foreground–background contrast.
Descriptive statistics (conditional PDFs)
Prediction Intersection Mean $P\left(z(x,y)\mid\text{Fg}\right)$ Mean $P\left(z(x,y)\mid\text{Bg}\right)$ Std $P\left(z(x,y)\mid\text{Fg}\right)$ Std $P\left(z(x,y)\mid\text{Bg}\right)$
Correct 0.24 0.57 0.16 0.14 0.22
Incorrect 0.32 0.33 0.14 0.01 0.21
Statistical validation (inside vs. outside mask)
Domain $z$-test (uncorrected) $z$-test Bonferroni $z$-test FDR (BH) Cohen’s $d$ (mean) Effect size interpretation
Energy ($z^{2}$) (98.2 %) (96.5 %) (98.2 %) 1.972 Very large effect
@@ 3.2.2 Expert-Model Agreement Assessment
Figure 4: Comprehensive example of the specialist segmentation process for pattern-based clinical validation. (a) shows the original dermoscopic image with multiple BCC patterns present; (b) displays the complete integrated segmentation combining all clinically relevant patterns identified by the dermatologist; (c)–(f) show individual pattern segmentations for Maple Leaf-like appearance (two instances), Ovoid Nests, and Multiglobules. This systematic segmentation approach enables pixel-level validation of the model’s attention mechanisms against clinically defined regions, ensuring that the AI tool’s visual explanations align with expert dermatological knowledge. The red contours indicate manually segmented regions of clinical interest used for quantitative comparison with Grad-CAM activation maps.
The conditional-PDF analysis shows that correctly classified lesions exhibit significantly higher activation inside clinically relevant regions than outside, reinforcing Grad-CAM as a trustworthy, clinically aligned explanation tool for BCC detection.
3.3 Model Selection and Deployment Considerations
A critical consideration in developing AI tools for clinical deployment is the balance between model performance and computational efficiency. In the context of teledermatology for primary care settings in Andalusia, where hardware resources are often limited and real-time inference is essential, we selected MobileNet-V2 as the backbone architecture. To validate this architectural choice, we compared its performance against more complex state-of-the-art architectures, specifically ConvNext Tiny and Vision Transformer (ViT-Base B/16), using the same training protocol and dataset.
Table 5 presents a comparative analysis of these architectures. While ConvNext Tiny and ViT-Base contain 8x and 25x more parameters respectively, and require 15x and 59x more computational operations (FLOPs), they do not provide meaningful improvements in pattern detection performance—the most challenging aspect of our task given the limited samples for minority BCC patterns. In fact, MobileNet-V2 achieves comparable or superior pattern detection accuracy (0.82) compared to ConvNext Tiny (0.77) and ViT-Base (0.78), despite being significantly more lightweight.
Table 5: Comparison of model architectures for BCC pattern detection and computational efficiency.
Model Parameters FLOPs BCC/Non-BCC Acc. Pattern Avg Acc.
MobileNet-V2 (ours) 3.5M 0.3G 0.90 0.82
ConvNext Tiny 28M 4.5G 0.90 0.77
ViT-Base (B/16) 86M 17.6G 0.86 0.78
This finding is particularly significant for clinical deployment: MobileNet-V2’s compact size (14 MB) and minimal computational requirements enable rapid inference on standard clinical workstations without requiring specialized GPU hardware, making it practical for integration into primary care workflows. The ability to deploy an effective diagnostic aid system without substantial infrastructure investment is essential for widespread adoption in resource-constrained healthcare settings. Our results demonstrate that architectural efficiency and clinical deployability need not come at the cost of diagnostic performance, particularly when the model is designed with domain-specific considerations such as pattern-based reasoning and clinical validation.
@@ 4 Discussion
Understanding the decision-making process of convolutional neural networks (CNNs) is critical for clinical applications where interpretability is as important as performance metrics. The most common visualization techniques in XAI are activation maps and gradient-weighted class activation mapping (Grad-CAM). Activation Maps, as described by Zhou et al. [zhou2015learning_CAM], extract features from different network layers to show which patterns the model focuses on, although they may not always directly correlate with specific outcomes. On the other hand, Grad-CAM, as described in detail by Selvaraju et al. [Selvaraju_2019_GradCAM], uses the gradients of the predicted class that flow into the final convolutional layer to produce a localization map. The clinically-inspired explanation that we proposed in this paper is based on Grad-CAM.
While gradient-based visualization methods such as Grad-CAM [Selvaraju_2019_GradCAM] are widely used in medical AI for post-hoc interpretability, our approach extends beyond standard visual explanations by integrating rigorous clinical validation through expert segmentations and conditional probability density functions. This quantitative validation ensures that the model’s attention aligns with diagnostically relevant regions identified by dermatologists, transforming Grad-CAM from a purely visual tool into a clinically validated explanation mechanism. Combined with our pattern-based reasoning layer, this dual-explanation framework provides both what the model detects (specific BCC patterns) and where it focuses (validated against expert knowledge), addressing the critical need for clinically aligned XAI in dermatological diagnosis.
@@
5 Conclusions and Implications
This work presents the first clinically validated dual-explanation AI system for Basal Cell Carcinoma detection that integrates pattern-based clinical reasoning directly into the model architecture. By combining Multi-Task Learning with explainable AI techniques, we bridge the critical gap between AI performance and clinical trust in teledermatology applications.
Our system achieves 90% accuracy in binary BCC/non-BCC classification while simultaneously providing clinically meaningful explanations through two complementary layers. The pattern-based explanation layer correctly identifies at least one BCC pattern in 99% of BCC-positive cases and properly recognizes the pigment network negative criterion in 95% of non-BCC cases. The visual explanation layer, validated through Grad-CAM analysis, demonstrates strong spatial agreement with dermatologist-segmented regions, with mean activation values of 0.57 inside clinically relevant areas compared to 0.16 in background regions. This dual-layer approach ensures that the AI system not only provides accurate diagnoses but also explanations that align with dermatological diagnostic criteria, addressing the interpretability requirements essential for clinical acceptance.
The clinical implications of this work are substantial. The lightweight MobileNet-V2 architecture enables efficient deployment in resource-constrained settings, making it particularly suitable for the teledermatology workflow between primary care centers and hospital dermatology units in Andalusia. By providing both a BCC/non-BCC diagnosis and pattern-level explanations, the system allows dermatologists to validate AI predictions through familiar clinical criteria, potentially reducing diagnostic errors and increasing confidence in AI-assisted decision-making. This approach directly addresses the current overload faced by hospital dermatologists managing cases from 60 primary care centers, offering a tool that not only aids in triage but also enhances the educational value of each case through transparent, clinically grounded explanations.
The integration of clinical reasoning into the model architecture, rather than relying solely on post-hoc explanation methods, represents a significant methodological advancement. Unlike previous approaches that use XAI techniques primarily to visualize regions of model attention, our MTL framework jointly learns classification and pattern detection tasks, enabling the system to reason about BCC in a manner aligned with clinical practice. This design ensures that explanations are not merely illustrative but are intrinsically connected to the diagnostic process itself.
Several avenues for future research emerge from this work. First, conducting reader studies with dermatologists would provide valuable insights into how the dual-explanation system affects diagnostic confidence and decision-making in real clinical scenarios. Second, extending the approach to other skin lesion types and incorporating additional dermoscopic criteria could broaden the system’s clinical utility. Third, investigating the integration of this tool into existing teledermatology platforms and evaluating its impact on workflow efficiency and diagnostic outcomes in prospective clinical trials would provide evidence for wider deployment. Finally, exploring the generalizability of the MTL-XAI framework to other medical imaging domains where clinical criteria guide diagnosis could extend the impact of this work beyond dermatology.
In conclusion, this work demonstrates that high-performance AI systems for medical diagnosis can be designed to provide transparent, clinically aligned explanations without sacrificing accuracy. By integrating pattern-based reasoning and visual explanations validated against expert annotations, we have developed a tool positioned for immediate deployment in primary care teledermatology, with the potential to improve both diagnostic efficiency and clinical trust in AI-assisted healthcare.